# The dusty envelopes of asymptotic giant branch stars with ultraviolet excesses

### 20 Mar 2026 v1

@@ 2.3 Herschel/PACS imaging
The relative flux uncertainties from the individual and combined Herschel/PACS images can be lower than . However, the flux variability between different epochs, even considering the scarcity of multi-epoch images and the sparse temporal coverage for most sources, is significantly larger (in some cases at least 5-20%). Therefore, the instrumental uncertainties underestimate the intrinsic variability of AGB stars on the far-infrared (this also can apply to IRAS and AKARI photometry). The employed Herschel/PACS observations and their respective estimated fluxes are summarised in appendix C.
@@ 3.1 Description of the DUSTY envelope models
Finally, the radial optical depth, , at a fiducial wavelength of 550 nm, is used to parametrise the optical depth of the CSE across the entire wavelength range covered in the model. We used the optical depths at the central wavelengths of the two GALEX bands in order to estimate the effect of dust attenuation in the observed UV emission (see appendix D).
@@ 3.2 Parameter space and SED fitting
Moreover, we noted that a few sources (R LMi, R UMa and Z Cnc) show a large difference between the fluxes obtained from Herschel/PACS and those from IRAS and AKARI (see Fig. 1). We noted that, in these cases, IRAS and AKARI fluxes are systematically higher, indicating pollution from an irregular sky brightness due to their more limited angular resolution. This large difference in a short wavelength range produces a large increase of and complicates the overall reliability of the fitting. Therefore, we did not include these fluxes in the analysis.
@@ 4.1 Best-fit solutions and uncertainties
We find that the SEDs of most of our targets (21/29) show an infrared excess and require a dusty CSE (see Figs. 1 and table 1); we derive additional parameters for these sources in Sects. 4.2 and 4.3. The SEDs of some sources (8/29) can be fitted using only stellar photospheric models (see Fig. 2 and table 2 implying a negligible dust component, presumably because they have not yet experienced substantial mass-loss. We refer to such sources as “naked” (as defined by Sloan & Price, 1995). Furthermore, we note that all the sources detected in CO by Alonso-Hernández et al. (2024) have infrared excess, whereas not all dusty AGB stars were detected in CO.
@@
Most of the sources presented in this study have optically-thin envelopes at the fiducial wavelength 1, although a few are optically-thick (1, see Fig.3 right). This clustering at low can be expected because uvAGBs with low optical depths have a higher chance of being detected via their UV emission due to a low attenuation (see Sect. 5 and Appendix D).
@@ 4.2 Derivation of related parameters
We derived the bolometric luminosity, the inner and outer radii of the CSE (, ), the radial optical depth as a function of wavelength (see also Appendix D), and the dust mass () from the spatial integration of the density distribution (see eqn. of Sahai et al., 2023).
@@ 4.4 Trends and comparison with larger samples
For this purpose, we performed a comparison with one independent stellar property and one extrinsic to the star. For the intrinsic stellar property, we used the variability classification and pulsation period () from Samus’ et al. (2017), summarised for this sample in Alonso-Hernández et al. (2024). For the extrinsic parameter, we used , which is associated with binarity (when 0.06, see Sahai et al., 2022) and large UV excesses (see Appendix. D).
@@
Fig. 5 shows the comparisons between the most relevant parameters estimated from the SED modelling with and (defined as corrected from dust attenuation, see appendix D). We evaluate these comparisons qualitatively because the large dependence between these parameters and different methodologies employed in the literature prevents the derivation of simple relationships to compare with (see discussion in Sect. 2 of Höfner & Olofsson, 2018)
@@ 5 Discussion
We also remark that binarity is expected to be common in AGB stars. In particular, the fraction of uvAGBs is at least 10% of the total population (see Alonso-Hernández et al., 2024). Therefore, it is likely that the uvAGBs present in the NESS sample (see Appendix D) are producing an overlap when comparing the samples, avoiding a clear identification of differences.
We also show that modelling dust attenuation at UV wavelengths is necessary for extracting intrinsic UV emission properties (see appendix. D). Optically thick CSEs can prevent even the detection of sources with intrinsically intense UV emission, resulting in an underestimate in the percentages of uvAGBs (as suggested by Montez et al., 2017).
We also highlight the effect of dust attenuation on the 0.06 threshold proposed by Sahai et al. (2022) as a criteria to discern cases in which the UV excesses are clearly related with accretion processes. The observed may well differ from the intrinsic ratio due to the difference in the optical depths between both bands. The optical depth in the FUV band is larger than in the NUV band, leading to intrinsic larger than observed. O-rich AGB stars present a larger difference in the optical depth between both bands (see appendix. D).
@@ 6 Summary & conclusions
The table containing the SED photometry will be available at the CDS.
@@ Acknowledgements.
We thank the referee, Roberto Ortiz, whose valuable comments helped us improve the quality of the manuscript. This work is part of the I+D+i projects PID2019-105203GB-C22, PID2022-137241NB-C42 and PID2023-146056NB-C22 funded by Spanish MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. J.A.H. is supported by INTA grant PRE_MDM_05 and acknowledges CSIC grant iMOVE 23023. R.S.’s contribution to the research described here was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA, and funded in part by NASA via various ROSES/ADAP awards and HST/GO awards (administered by STScI). Herschel is an ESA space observatory with science instruments provided by European-led Principal Investigator consortia and with important participation from NASA. ISO is an ESA project with instruments funded by ESA Member States and with the participation of ISAS and NASA. Some of the data presented in this paper were obtained from the Multimission Archive at the Space Telescope Science Institute (MAST). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555. Support for MAST for non-HST data is provided by the NASA Office of Space Science via grant NAG5-7584 and by other grants and contracts.References
@@ Appendix B VY UMa detached shell
Appendix C Summary of Herschel/PACS observations
We present the compilation of photometric measurements for all the individual Herschel/PACS observation employed in this study. Table 3 shows the technical information and the individual measurements obtained from aperture photometry in the two employed bands of each observation.
[b] Source OBs ID OBs date duration mode Flux band B Flux band G Flux band R (MJD) (s) (Jy) (Jy) (Jy) AT Dra 1342190578 55239.274 152 chop 1.190.03 – 0.2080.009 1342191845 55264.457 152 chop 1.230.03 – 0.2580.008 EY Hya 1342196885 55338.979 26960 parallel – 2.510.29 0.940.09 1342196886 55339.293 24690 parallel – 2.510.29 0.940.09 1342197282 55345.144 26960 parallel – 1.820.21 0.750.10 1342197283 55345.457 24690 parallel – 1.820.21 0.750.10 1342253344 56211.138 52 scan map 4.050.01 – 0.730.03 1342253500 56215.995 52 scan map 4.140.02 – 0.970.17 IN Hya 1342197307 55345.979 26960 parallel – 0.250.06 0.13 1342197308 55346.292 26960 parallel – 0.250.06 0.13 ome Vir 1342187362 55165.745 161 chop-nod 1.510.03 – 0.260.06 1342198902 55368.940 152 chop-nod 1.480.03 – 0.290.06 1342223317 55741.738 152 chop-nod 1.480.02 – 0.330.05 1342223625 55745.669 152 chop-nod 1.450.02 – 0.270.05 1342223690 55746.726 152 chop-nod 1.510.02 – 0.310.04 1342223885 55752.614 152 chop-nod 1.510.03 – 0.260.05 1342223925 55753.611 152 chop-nod 1.580.03 – 0.260.06 R LMi 1342254139 56230.736 52 scan map 11.720.02 – 2.450.05 R UMa 1342221940 55712.986 52 scan map 2.810.05 – 0.470.05 1342246162 56073.792 52 scan map 2.990.04 – 0.560.04 1342246187 56074.789 52 scan map 3.120.06 – 0.590.05 1342246747 56085.991 52 scan map 3.270.06 – 0.690.04 1342246935 56088.951 52 scan map 3.330.06 – 0.570.02 RR Eri 1342190602 55239.380 152 chop-nod 1.750.02 – 0.350.05 RR UMi 1342254046 56229.690 52 scan map 2.970.01 – 0.530.02 RT Cnc 1342253343 56211.136 52 scan map 2.300.07 – 0.430.04 1342253469 56215.813 52 scan map 2.440.06 – 0.560.06 RU Her 1342190522 55239.065 152 chop-nod 8.370.03 – 1.430.03 1342191825 55264.384 152 chop-nod 8.220.03 – 1.380.04 1342228878 55822.857 152 chop-nod 6.610.03 – 1.140.03 ST UMa 1342247296 56101.709 52 scan map 1.630.03 – 0.320.09 1342247425 56103.683 52 scan map 1.630.04 – 0.190.06 SV Peg 1342220943 55696.010 152 chop-nod 13.030.05 – 1.870.16 1342221173 55697.791 152 chop-nod 13.170.04 – 1.830.12 1342222726 55729.050 152 chop-nod 12.880.05 – 1.840.11 • Notes. Column (1): Source, Col. (2): Obs ID, Col. (3): OBs date, Col. (4): exposure duration, Col. (5): observational mode, Col. (6): Flux in band B, Col. (7): Flux in band G, Col. (8): Flux in band R. Herschel/PACS B, G and R filters have effective wavelengths of 72, 103 and 167 m, and effective spectral widths of 21, 31, and 70 m respectively.
[b] Source OBs ID OBs date duration mode Flux band B Flux band G Flux band R (MJD) (s) (Jy) (Jy) (Jy) T Dra 1342187357 55165.608 161 chop-nod 8.470.02 – 1.660.03 1342190576 55239.267 152 chop-nod 7.760.02 – 1.550.03 1342222753 55729.353 152 chop-nod 8.660.02 – 1.620.03 1342232432 55882.643 152 chop-nod 10.000.02 – 2.080.03 1342232465 55883.675 152 chop-nod 11.100.02 – 2.020.03 1342233065 55891.711 152 chop-nod 11.130.02 – 2.020.03 1342233205 55892.709 152 chop-nod 11.020.02 – 2.030.03 1342233356 55896.707 152 chop-nod 11.040.02 – 1.990.03 1342233587 55899.624 152 chop-nod 11.070.02 – 1.980.03 1342234216 55900.630 152 chop-nod 10.930.02 – 2.000.03 1342240701 55992.767 152 chop-nod 9.090.02 – 1.690.03 1342241298 55999.595 152 chop-nod 8.930.02 – 1740.03 V Eri 1342189943 55227.184 283 scan map 12.610.04 – 1.900.16 1342189944 55227.188 283 scan map 12.610.04 – 1.900.16 VY UMa 1342210606 55530.947 881 scan map 1.200.04 – 0.190.02 1342210607 55530.958 881 scan map 1.200.04 – 0.190.02 1342243219 56008.669 733 scan map – 0.640.02 0.230.03 1342243220 56008.678 733 scan map – 0.640.02 0.230.03 W Peg 1342222729 55729.060 152 chop-nod 6.610.03 – 1.280.04 Z Cnc 1342251911 56200.817 52 scan map 1.330.01 – 0.200.13
Appendix D UV excesses and dust attenuation
@@
We computed synthetic stellar photometry on the GALEX bands and compared them with the observed fluxes to estimate the UV excesses (similarly to Sahai et al. 2008; Ortiz & Guerrero 2016). We noted that GALEX NUV and FUV photometric bands fall outside the wavelength range covered by the COMARCS models used in this work (see Sect. 3). Therefore, the stellar fluxes were estimated by extrapolating the input spectra following a blackbody distribution and convolving it with the transmission curves of their respective filters. We acknowledge that the real stellar spectrum differs from a blackbody approximation mainly due to the presence of (i) molecular absorption bands, and (ii) emission-line features due to possible chromospheric emission, that can produce a significant effect on the shape of the stellar spectra in the UV. The features in (ii) are expected to have a much stronger effect in the FUV (compared to the NUV) where the photospheric contribution is relatively much weaker. Note that the sole purpose of this simple approximation is to explore whether is a reasonable proxy for the FUV excess.
@@ Appendix F Comparison with NESS sample
Fig. 16 shows the comparison between the gas-to-dust ratio and the ratio between CO and IRAS 60 m and . It is shown that correlates with the ratio between CO and infrared emission, although some outliers (likely related with pollution in the IRAS 60 m images) are present. We noted that is not a empirical value and depends on the employed assumptions and methodology. Therefore, it is likely the presence of some systematic differences between our estimates and NESS. On the other hand, anticorrelates with , this correlation, although it presents some scatter, which can be related with UV emission variability, and it is not corrected from dust attenuation.

### 27 Mar 2026 v2

@@ 2.3 Herschel/PACS imaging
The relative flux uncertainties from the individual and combined Herschel/PACS images can be lower than . However, the flux variability between different epochs, even considering the scarcity of multi-epoch images and the sparse temporal coverage for most sources, is significantly larger (in some cases at least 5-20%). Therefore, the instrumental uncertainties underestimate the intrinsic variability of AGB stars on the far-infrared (this also can apply to IRAS and AKARI photometry). The employed Herschel/PACS observations and their respective estimated fluxes are summarised in electronic format at the CDS.
@@ 3.1 Description of the DUSTY envelope models
Finally, the radial optical depth, , at a fiducial wavelength of 550 nm, is used to parametrise the optical depth of the CSE across the entire wavelength range covered in the model. We used the optical depths at the central wavelengths of the two GALEX bands in order to estimate the effect of dust attenuation in the observed UV emission (see appendix C).
@@ 3.2 Parameter space and SED fitting
Moreover, we noted that a few sources (R LMi, R UMa and Z Cnc) show a large difference between the fluxes obtained from Herschel/PACS and those from IRAS and AKARI (see Fig. 1 and appendix D). We noted that, in these cases, IRAS and AKARI fluxes are systematically higher, indicating pollution from an irregular sky brightness due to their more limited angular resolution. This large difference in a short wavelength range produces a large increase of and complicates the overall reliability of the fitting. Therefore, we did not include these fluxes in the analysis.
@@ 4.1 Best-fit solutions and uncertainties
We find that the SEDs of most of our targets (21/29) show an infrared excess and require a dusty CSE (see Fig. 1, table 1 and appendix D); we derive additional parameters for these sources in Sects. 4.2 and 4.3. The SEDs of some sources (8/29) can be fitted using only stellar photospheric models (see Fig. 2 and table 2 implying a negligible dust component, presumably because they have not yet experienced substantial mass-loss. We refer to such sources as “naked” (as defined by Sloan & Price, 1995). Furthermore, we note that all the sources detected in CO by Alonso-Hernández et al. (2024) have infrared excess, whereas not all dusty AGB stars were detected in CO.
@@
Most of the sources presented in this study have optically-thin envelopes at the fiducial wavelength 1, although a few are optically-thick (1, see Fig.3 right). This clustering at low can be expected because uvAGBs with low optical depths have a higher chance of being detected via their UV emission due to a low attenuation (see Sect. 5 and Appendix C).
@@ 4.2 Derivation of related parameters
We derived the bolometric luminosity, the inner and outer radii of the CSE (, ), the radial optical depth as a function of wavelength (see also Appendix C), and the dust mass () from the spatial integration of the density distribution (see eqn. 1 of Sahai et al., 2023).
@@ 4.4 Trends and comparison with larger samples
For this purpose, we performed a comparison with one independent stellar property and one extrinsic to the star. For the intrinsic stellar property, we used the variability classification and pulsation period () from Samus’ et al. (2017), summarised for this sample in Alonso-Hernández et al. (2024). For the extrinsic parameter, we used , which is associated with binarity (when 0.06, see Sahai et al., 2022) and large UV excesses (see Appendix C).
@@
Fig. 5 shows the comparisons between the most relevant parameters estimated from the SED modelling with and (defined as corrected from dust attenuation, see appendix C). We evaluate these comparisons qualitatively because the large dependence between these parameters and different methodologies employed in the literature prevents the derivation of simple relationships to compare with (see discussion in Sect. 2 of Höfner & Olofsson, 2018)
@@ 5 Discussion
We also remark that binarity is expected to be common in AGB stars. In particular, the fraction of uvAGBs is at least 10% of the total population (see Alonso-Hernández et al., 2024). Therefore, it is likely that the uvAGBs present in the NESS sample (see Appendix C) are producing an overlap when comparing the samples, avoiding a clear identification of differences.
We also show that modelling dust attenuation at UV wavelengths is necessary for extracting intrinsic UV emission properties (see appendix C). Optically thick CSEs can prevent even the detection of sources with intrinsically intense UV emission, resulting in an underestimate in the percentages of uvAGBs (as suggested by Montez et al., 2017).
We also highlight the effect of dust attenuation on the 0.06 threshold proposed by Sahai et al. (2022) as a criteria to discern cases in which the UV excesses are clearly related with accretion processes. The observed may well differ from the intrinsic ratio due to the difference in the optical depths between both bands. The optical depth in the FUV band is larger than in the NUV band, leading to intrinsic larger than observed. O-rich AGB stars present a larger difference in the optical depth between both bands (see appendix C).
@@ 6 Summary and conclusions
Tables containing the employed photometry for the SEDs and the photometric fluxes estimated from the individual Herschel/PACS observations are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
We thank the referee, Roberto Ortiz, whose valuable comments helped us improve the quality of the manuscript. This work is part of the I+D+i projects PID2019-105203GB-C22, PID2022-137241NB-C42 and PID2023-146056NB-C22 funded by Spanish MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. J.A.H. is supported by INTA grant PRE_MDM_05 and acknowledges CSIC grant iMOVE 23023. R.S.’s contribution to the research described here was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA (80NM0018D0004), and funded in part by NASA via various ROSES/ADAP awards and HST/GO awards (administered by STScI). Herschel is an ESA space observatory with science instruments provided by European-led Principal Investigator consortia and with important participation from NASA. ISO is an ESA project with instruments funded by ESA Member States and with the participation of ISAS and NASA. Some of the data presented in this paper were obtained from the Multimission Archive at the Space Telescope Science Institute (MAST). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555. Support for MAST for non-HST data is provided by the NASA Office of Space Science via grant NAG5-7584 and by other grants and contracts.References
@@ Appendix B VY UMa detached shell
Appendix C UV excesses and dust attenuation
@@
We computed synthetic stellar photometry on the GALEX bands and compared them with the observed fluxes to estimate the UV excesses (similarly to Sahai et al. 2008; Ortiz & Guerrero 2016). We noted that GALEX NUV and FUV photometric bands fall outside the wavelength range covered by the COMARCS models used in this work (see Sect. 3). Therefore, the stellar fluxes were estimated by extrapolating the input spectra following a blackbody distribution and convolving it with the transmission curves of their respective filters. Table 3 shows the observed and modelled properties of the UV emission for the sources used in this sample.
We acknowledge that the real stellar spectrum differs from a blackbody approximation mainly due to the presence of (i) molecular absorption bands, and (ii) emission-line features due to possible chromospheric emission, that can produce a significant effect on the shape of the stellar spectra in the UV. The features in (ii) are expected to have a much stronger effect in the FUV (compared to the NUV) where the photospheric contribution is relatively much weaker. Note that the sole purpose of this simple approximation is to explore whether is a reasonable proxy for the FUV excess.
@@
Appendix D SEDs and best-fit models for the rest of dusty AGB stars
Fig. 14 shows the SEDs and their best-fit DUSTY models for the rest of dusty AGB stars to complement the sources already shown in Fig. 1.
@@ Appendix F Comparison with NESS sample
Fig. 17 shows the comparison between the gas-to-dust ratio and the ratio between CO and IRAS 60 m and . It is shown that correlates with the ratio between CO and infrared emission, although some outliers (likely related with pollution in the IRAS 60 m images) are present. We noted that is not a empirical value and depends on the employed assumptions and methodology. Therefore, it is likely the presence of some systematic differences between our estimates and NESS. On the other hand, anticorrelates with , although it presents some scatter that can be related with UV emission variability and it is not corrected from dust attenuation.